# Cellular Wireless Networks in the Upper Mid-Band

### 11 Dec 2023 v5

@@ Abstract
The upper mid-band – roughly from 7 to 24 — has attracted considerable recent interest for new cellular services. This frequency range has vastly more spectrum than the highly congested bands below while offering more favorable propagation and coverage than the millimeter wave (mmWave) frequencies. The upper mid-band can thus provide a powerful and complementary frequency range to balance coverage and capacity. Realizing the full potential of these bands, however, will require fundamental changes to the design of cellular systems. Most importantly, spectrum will likely need to be shared with incumbents including communication satellites, military RADAR, and radio astronomy. Also, the upper mid-band is simply a vast frequency range. Due to this wide bandwidth, combined with the directional nature of transmission and intermittent occupancy of incumbents, cellular systems will need to be agile to sense and intelligently use large spatial and frequency degrees of freedom. This paper attempts to provide an initial assessment of the feasibility and potential gains of wideband cellular systems operating in the upper mid-band. The study includes: (1) a system study to assess potential gains of multi-band systems in a representative dense urban environment and illustrate the value of wide band system with dynamic frequency selectivity; (2) an evaluation of potential cross interference between satellites and terrestrial cellular services and interference nulling to reduce that interference; and (3) design and evaluation of a compact multi-band antenna array structure. Leveraging these preliminary results, we identify potential future research directions to realize next-generation systems in these frequencies.
Index Terms:
Upper mid-band, 6G, cellular wireless systems, FR3, satelliteI Introduction
@@
Cellular wireless systems up to the 4G had largely operated in a range of microwave frequencies below . Given the severe spectral shortage in these bands, 5G systems [dahlman20205g, 3GPP38300] introduced new capabilities in the millimeter wave (mmWave) frequencies above [RanRapE:14, AkdenizCapacity:14, shafi20175g]. In the US, individual carriers had spectrum allocations as large as – several times larger than previous allocations in the sub-6 bands. These wide bandwidths, combined with spatial multiplexing capabilities, enabled massive multi-Gbps peak rates. However, extensive measurements have now shown that practical performance is often intermittent [narayanan2021variegated, wei20225gperf] with limited penetration inside [narayanan2022comparative, rochman2022comparison]. This poor coverage is due to the inherently limited range of mmWave signals and their high susceptibility to blockage [Rappaport2014-mmwbook, slezak2018empirical, slezak2018understanding, maccartney2017rapid, zekri2020analysis].
@@
Against this backdrop, the upper mid-band spectrum, roughly from 7-24 , provides a good balance of coverage and spectrum, overcoming the spectral shortage of the sub-6 bands while having favorable propagation and penetration relative to the mmWave bands. For these reasons, the band has recently attracted considerable attention for commercial cellular communications and has been cited by industry as a leading candidate spectrum for next generation (NextG, 5G and beyond) wireless networks [smee2022ten, samsung20226gspectrum, 5Gamerica20223gpproadmap]. The Federal Communications Commission (FCC) Technical Advisory Committee (TAC) has recently identified the upper mid-band as vital for meeting the growing data rate demands [fcc2023preliminary, fcc20236Gworkinggroup]. The 3rd Generation Partnership Project (3GPP), the organization that sets cellular standards, has also begun study of these bands [3gpp38820] – see Section II for more details.
@@
In 3GPP, the upper mid-band is referred to as frequency range 3 (FR3). FR1 initially referred to the traditional spectrum below , which was later expanded to [sathya2020standardization]. FR2 refers to the mmWave spectrum that was introduced in the fifth generation (5G) standard).
I-A Challenges
@@
Need for spectrum sharing and sensing with incumbents: Most importantly, the bands are already used by several vital incumbent services. As we review in Section II-C, commercial satellite services already widely use these bands and have been interested in increased bandwidth allocations, particularly for broadband rural access. The upper mid-band also contains scientifically important spectral lines that are fundamental for radio astronomy as well as vital military RADAR.
Large spectral range: The upper mid-band is simply a vast range of spectrum – well beyond the span of most commercial cellular front-ends that typically operate over a small percentage of the carrier frequencies. This large range, combined with the need to dynamically share spectrum and directionally transmit, demands new front-ends that adaptively and intelligently sense and exploit large bandwidth and spatial degrees of freedom.
@@
I-B Contributions
The broad goal of this paper is to assess the feasibility and potential of agile wideband cellular systems operating in these bands. We will also identify key challenges and areas for research to realize next-generation systems in these bands. Towards this end, our contributions are:
Channel modeling and estimation of multi-band coverage: We briefly review channel measurements in the upper mid-band range and discuss open areas of research (Section III-A). Also, based on current propagation models, we conduct a ray-tracing simulation in a representative dense urban area (New York City) to assess the coverage and capacity of a potential cellular system in the upper mid-band (Section III-B). Our results demonstrate that mobiles that can adaptively select one of multiple bands across the upper mid-band can improve coverage and data rates over mobiles restricted to a single band. This finding argues for multi-band, adaptive systems to gain the full benefits of the frequency range.
@@ I-B Contributions
II Background and Standardization Landscape
II-A Prior developments in the sub-6 and mmWave bands
The interest in the upper mid-band has to be seen in the context of developments of commercial cellular systems in both the sub-7 and mmWave frequency ranges over the last decade. Early mmWave experiments and capacity analyses such as [akoum2012coverage, rappaportmillimeter, RanRapE:14, AkdenizCapacity:14] suggested the possibility of massive data rates in micro-cellular deployments owing to the wide available bandwidths and spatial multiplexing gains available in the mmWave frequencies. Based partially on these and other results, the mmWave bands have emerged as an integral component of the 5G New Radio (NR) specification [dahlman20205g, 3GPP38300]. Commercial mmWave deployments appeared shortly after the release of the speceification, particularly in the US [shafi20175g].
Coverage at reasonable cell densities in these deployments, however, has been an enormous challenge. Several recent measurements in commercial mmWave networks now clearly demonstrate that outdoor coverage can be highly intermittent [narayanan2021variegated, wei20225gperf] with limited penetration indoors [narayanan2022comparative, rochman2022comparison]. MmWave signals are simply blocked by many common building materials such as concrete [Rappaport2014-mmwbook, zekri2020analysis] as well as the human body and other obstacles [slezak2018empirical, slezak2018understanding, maccartney2017rapid]. See also Section III-D below.
@@ II-A Prior developments in the sub-6 and mmWave bands
Measurements from commercial network monitoring companies, now offer an opportunity to compare the practical performance of these networks. On the one hand, Ookla reported in 2022 [ookla2023mmwaverates] that the tested mmWave networks offer an incredible median downlink throughput, up to seven times higher than systems in C-band. Similarly, in 2021, Open Signal demonstrated [opensignal2021mmwave] that Verizon’s mmWave downlink capacity was at least three times as high as networks relying mostly on mid-band spectrum. Nevertheless, of mid-band networks enabled downlink speeds in the hundreds of Mbps. Moreover, the coverage was much more uniform. For example, Open Signal’s study [opensignal2021mmwave] showed that users connect to mmWave less than 1% of the time.
II-B Recent standardization efforts in the upper mid-band
The relative success of commercial cellular systems in the mid-band, combined with the high data rates, but intermittent coverage, of mmWave has set the stage for interest in the upper mid-band. The hope is to provide the high data rates close to those available in the mmWave range, but with much more uniform coverage. As mentioned in the Introduction, the bands have been identified by industry [smee2022ten, samsung20226gspectrum] and 3GPP has recently formally started study of services in the 7 to 24 band [3gpp38820, 5Gamerica20223gpproadmap], where they are called frequency range 3 or FR3.
Note that, in 3GPP terminology, frequency range 1, or FR1, has been extended from sub-6 GHz to sub-7 GHz to include the frequencies in the 6 to 7 range [3GPP38921]. Nevertheless, many discussions use the term upper mid-band to refer to 6 to 24 .
In 2022, the FCC also began discussion on two bands from 12.2-12.7 [FCC20-443], and 12.7-13.2 [FCC22-352] for possible cellular use. While the first band has now been rejected for cellular services, there is wide recognition of the need for opening the upper mid-band for cellular services. For example, the FCC Technical Advisory Committee (TAC) has recently published a comprehensive survey of spectrum from 7.125 to 24 [fcc2023preliminary] as well as 6G working paper [fcc20236Gworkinggroup]. These analyses emphasize the need for considering larger portions of the upper mid-band for terrestrial cellular services in order to meet the growing data demand.
II-C Incumbency and the need for spectrum sharing
@@
III Coverage and Capacity Gains
III-A Prior channel measurements and studies
@@
III-B Multi-frequency outdoor capacity gains
@@
To estimate the potential gain of such a multi-band, frequency adaptive system in the upper mid-band, we consider a dense urban area, i.e., Herald Square in New York City, as shown in Fig. 2a and downlink scenario (gNBUE). In dense urban scenarios, capacity is mostly limited, while providing satisfactory coverage is challenging at high frequencies due to blockage, a key challenge, for example, that has been experienced in the mmWave bands [slezak2018empirical, slezak2018understanding, jain2018driven]. Within this area, we see systems operating at up to four potential frequencies: 6, 12, 18, and 24.
@@
The simulation parameters of the potential cellular downlink systems at these frequencies are shown in Table I. Note that the bandwidth in each frequency scales with the carrier, as is typical in deployments today. As shown in Fig. 2b, the size of the gNB antenna array also scales so that the aperture is approximately constant. Furthermore, in the simulation area, we manually selected the locations of 18 BSs (gNB) on rooftops, corresponding to an inter-site distance (ISD) of approximately , typical for urban microcellular evaluations. Terrestrial UEs are randomly placed outside buildings, and ray tracing was performed using Wireless Insite [Remcom], which has recently been proven fairly accurate by conducting real-world measurements in the mmWave bands, as reported in [de2023ray, remcommpresnt2017modeling].
We compute the wideband SNR for each UE associated with the serving BS. At each UE location, we select one BS serving with the strongest SNR. Fig. 2c shows a coverage map for a single BS deployment as an example. We see that coverage is greatly reduced as frequency increases. Similarly, the curves in Fig. 3 for frequencies 6, 9, 18 and 24 represent the cumulative distribution function (CDF) of the signal-to-noise ratio (SNR) and rate for a hypothetical UE.
@@
To compute the achieved rate (goodput) with consideration of system overheads, we use a standard realistic model [mogensen2007lte] as the follows
@@
where is the bandwidth, is a system bandwidth loss factor to account for overhead and receiver imperfections, and is a maximum spectral efficiency. As indicated in [mogensen2007lte], we adopt for and for .
@@
As shown in Fig. 3a, UEs at lower frequencies experience better SNRs due to benign propagation conditions and reduced noise power using a smaller bandwidth compared to higher frequencies. We plot the CDFs of the data rate for each frequency in Fig. 3b, where the maximum data rates for , , and are specified as , , and , respectively. As expected, the peak rates for UEs are higher at higher frequencies (18 and 24 ) due to wider bandwidths. However, cell edge UEs (for example, UEs up to the bottom 10% percentile of the CDF) suffer a significantly worse rate in the higher frequency bands (18 and 24 ) compared to UEs in the lower frequencies (6 and 12 ). This property is also expected, as propagation is much less favorable for UEs at the edge of the cell at higher frequencies.
@@
The curves labeled “best choice” in Fig. 3b correspond to the rates for the UEs that select the best BS and frequency. They achieve the same peak rates as UEs in the high-frequency bands. At the same time, we observe that the cell edge rate is improved over any individual frequency. These results suggest that cellular systems that adaptively select across a range of bands through the given spectrum are able to obtain peak rates at higher carrier frequencies while improving cell edge rates.
III-C Penetration Loss
@@
III-D Multi-frequency Indoor Capacity Gains
The 3D models used in the ray tracing have approximate material classifications for the facades of each building. Using the penetration loss models from Eq. 2 combined with the exterior wall classification, Fig. 5 plots the estimated CDFs of the SNR and data rate for indoor UEs. Fig. 5a shows that the SNR for indoor UEs is considerably reduced at higher frequencies due to high penetration loss, as mentioned in the previous section. In this setting, concrete is a dominant exterior wall material and significantly reduces the signal penetration – see Fig. 4. Fig. 5b shows that, for roughly 65% of the users, the data rates at lower frequencies are higher than at higher frequencies. These results indicate a further potential gain of wideband cellular systems in FR3: the lower frequencies can provide valuable indoor coverage behind materials such as concrete while the higher frequencies can opportunistically offer high capacity for outdoor users and selected indoor users with minimal blockage (e.g., indoor users next to non-IRR glass windows).
III-E Effects of Interference

### 06 Mar 2024 v6

# OJCOMS.2024.011100

@@ Abstract
The upper mid-band – approximately from 7 to 24 — has recently attracted considerable interest for new cellular services. This frequency range has vastly more spectrum than the highly congested bands below while offering more favorable propagation and coverage than the millimeter wave (mmWave) frequencies. In this regard, the upper mid-band has the potential to provide a powerful and complementary frequency range that balances coverage and capacity. Realizing cellular networks that exploit the full range of these bands, however, presents significant technical challenges. Most importantly, spectrum will likely need to be shared with incumbents including communication satellites, military RADAR, and radio astronomy. Also, the upper mid-band is simply a vast frequency range. Due to this wide bandwidth, combined with the directional nature of transmission and intermittent occupancy of incumbents, cellular systems will likely need to be agile to sense and intelligently use large spatial and frequency degrees of freedom. This paper attempts to provide an initial assessment of the feasibility and potential gains of such adaptive wideband cellular systems operating across the upper mid-band. The study includes: (1) a detailed ray tracing simulation to assess potential gains of multi-band systems in a representative dense urban environment and illustrate the value of a wideband system with dynamic frequency selectivity; (2) an evaluation of potential cross-interference between satellites and terrestrial cellular services and interference nulling to reduce that interference; and (3) design and evaluation of a compact multi-band antenna array structure. Leveraging these preliminary results, we identify potential future research directions to realize next-generation systems in these frequencies.
Upper mid-band, 6G, cellular wireless systems, FR3, satellite communications
1 Introduction
@@
Cellular wireless systems up to the fourth generation (4G) had largely operated in a range of microwave frequencies below . Given the severe spectral shortage in these bands, 5G systems [dahlman20205g, 3GPP38300] introduced new capabilities in the millimeter wave (mmWave) frequencies above [RanRapE:14, AkdenizCapacity:14, shafi20175g]. The wide bandwidths available in the mmWave range, combined with spatial multiplexing capabilities, have enabled massive multi-Gbps peak rates. However, extensive measurements also now show that practical performance is often intermittent [narayanan2021variegated, wei20225gperf] with limited penetration inside [narayanan2022comparative, rochman2022comparison]. At root, this poor coverage is due to the inherently limited range of mmWave signals and their high susceptibility to blockage [Rappaport2014-mmwbook, slezak2018empirical, slezak2018understanding, maccartney2017rapid, zekri2020analysis].
@@
Against this backdrop, the upper mid-band spectrum, roughly from 7-24 , is being considered to provide a good balance of coverage and bandwidth, overcoming the spectral shortage of the sub-6 bands while having favorable propagation and penetration relative to the mmWave frequencies. For these reasons, the band has recently attracted considerable attention for commercial cellular communications and has been cited by industry as a leading candidate spectrum for next generation (NextG, 5G and beyond) wireless networks [smee2022ten, samsung20226gspectrum, 5Gamerica20223gpproadmap, lee2022extreme]. The Federal Communications Commission (FCC) Technical Advisory Committee (TAC) has recently identified the upper mid-band as vital for meeting the growing data rate demands [fcc2023preliminary, fcc20236Gworkinggroup]. The 3rd Generation Partnership Project (3GPP), the organization that sets cellular standards, has also begun study of these bands [3gpp38820] – see Section 2 for more details.
@@
In 3GPP, the upper mid-band is referred to as Frequency Range 3 (FR3). FR1 initially referred to the traditional spectrum below , which was later expanded to [sathya2020standardization]. FR2 refers to the mmWave spectrum that was introduced in the fifth generation (5G) standard.
1.1 Challenges
@@
Need for sensing and spectrum sharing with incumbents: Most importantly, these bands are already used by several vital incumbent services. This situation stands in contrast to the mmWave deployments where the bands were significantly less occupied. As we review in Section 2.3, commercial satellite services already widely use these bands and have been interested in increased bandwidth allocations, particularly for broadband rural access. The upper mid-band is also vital for military and commercial RADAR and contains scientifically important spectral lines that are fundamental for radio astronomy.
Large spectral range: The upper mid-band is simply a vast range of spectrum – well beyond the span of most commercial cellular front-ends that typically operate over a small percentage of the carrier frequency. This large range, combined with the need to dynamically share spectrum and directionally transmit, demands new front-ends that adaptively and intelligently sense and exploit large bandwidth and spatial degrees of freedom.
@@
1.2 Contributions
The broad goal of this paper is to assess the feasibility and potential of agile wideband cellular systems operating in these bands. We will also identify key challenges and areas for research. Towards this end, our contributions are:
Channel modeling and estimation of multi-band coverage: We briefly review channel measurements in the upper mid-band range and discuss open areas of research (Section 3.1). Also, based on current propagation models, we conduct a ray-tracing simulation in a representative dense urban area (New York City) to assess the coverage and capacity of a potential cellular system in the upper mid-band (Section 3.2). Our results demonstrate that mobiles that can adaptively select one of multiple bands across the upper mid-band can improve coverage and data rates over mobiles restricted to a single band. This finding argues for multi-band, adaptive systems to gain the full benefits of the frequency range.
@@ 1.2 Contributions
2 Background and Standardization Landscape
2.1 Prior Developments in the Sub-6 and mmWave Bands
The interest in the upper mid-band has to be seen in the context of developments of commercial cellular systems in both the sub-7 and mmWave frequency ranges over the last decade. Early mmWave experiments and capacity analyses such as [akoum2012coverage, rappaportmillimeter, RanRapE:14, AkdenizCapacity:14] suggested the possibility of massive data rates in micro-cellular deployments owing to the wide bandwidths and spatial multiplexing gains available in the mmWave frequencies. Based partially on these and other results, the mmWave bands emerged as an integral component of the 5G New Radio (NR) specification [dahlman20205g, 3GPP38300]. Commercial mmWave deployments appeared shortly after the release of the specification, particularly in the US [shafi20175g].
Coverage at reasonable cell densities in these deployments, however, has been an enormous challenge. Several recent measurements in commercial mmWave networks now clearly demonstrate that outdoor coverage can be highly intermittent [narayanan2021variegated, wei20225gperf] with limited penetration indoors [narayanan2022comparative, rochman2022comparison]. MmWave signals are simply blocked by many common building materials such as concrete [Rappaport2014-mmwbook, zekri2020analysis] as well as the human body and other obstacles [slezak2018empirical, slezak2018understanding, maccartney2017rapid]. See also Section 3.4 below.
@@ 2.1 Prior Developments in the Sub-6 and mmWave Bands
Measurements from commercial network monitoring companies, now offer an opportunity to compare the practical performance of these networks. On the one hand, Ookla reported in 2022 [ookla2023mmwaverates] that the tested mmWave networks offer an incredible median downlink throughput, up to seven times higher than systems in C-band. Similarly, in 2021, Open Signal demonstrated [opensignal2021mmwave] that Verizon’s mmWave downlink capacity was at least three times as high as networks relying mostly on mid-band spectrum. Nevertheless, networks with mid-band alone often provided downlink speeds in the hundreds of Mbps. Moreover, the coverage for mid-band networks was much more uniform than mmWave. For example, Open Signal’s study [opensignal2021mmwave] showed that users connect to mmWave less than 1% of the time.
2.2 Recent Standardization Efforts in the Upper Mid-band
The relative success of commercial cellular systems in the mid-band, combined with the high data rates, but intermittent coverage, of mmWave has set the stage for interest in the upper mid-band. The hope is to provide the high data rates close to those available in the mmWave range, but with much more uniform coverage. As mentioned in the Introduction, the bands have been identified by industry [smee2022ten, samsung20226gspectrum] and 3GPP has recently formally started study of services in the 7 to 24 band [3gpp38820, 5Gamerica20223gpproadmap], where they are called frequency range 3 or FR3. Note that, in 3GPP terminology, frequency range 1, or FR1, has been extended from sub-6 GHz to sub-7 GHz to include the frequencies in the 6 to 7 range in FR1 [3GPP38921].
In 2022, the FCC also began discussion on two bands from 12.2-12.7 [FCC20-443], and 12.7-13.2 [FCC22-352] for possible cellular use. While the first band has now been rejected for cellular services, there is wide recognition of the need for opening the upper mid-band for cellular services. For example, the FCC Technical Advisory Committee (TAC) has recently published a comprehensive survey of spectrum from 7.125 to 24 [fcc2023preliminary] as well as a 6G working paper [fcc20236Gworkinggroup]. These analyses emphasize the need for considering larger portions of the upper mid-band for terrestrial cellular services in order to meet the growing data demand.
2.3 Incumbency and the Need for Spectrum Sharing
@@
3 Coverage and Capacity Gains
3.1 Prior Channel Measurements and Studies
@@
3.2 Multi-frequency Outdoor Capacity Gains
@@
To estimate the potential gain of such a multi-band, frequency adaptive system in the upper mid-band, we consider a dense urban area, i.e., Herald Square in New York City, as shown in Fig. 2a and downlink scenario (gNBUE). In dense urban scenarios, capacity a key requirement. At the same time, providing satisfactory coverage is challenging at high frequencies due to blockage, as have been experienced in the mmWave bands [slezak2018empirical, slezak2018understanding, jain2018driven]. Within this area, we consider a hypothetical systems operating at up to four potential frequencies: 6, 12, 18, and 24.
@@
The simulation parameters of the potential cellular downlink systems at these frequencies are shown in Table 1. Note that the bandwidth in each frequency scales with the carrier, as is typical in deployments today. As shown in Fig. 2b, the size of the gNB antenna array also scales so that the aperture is approximately constant. Furthermore, in the simulation area, we manually selected the locations of 18 BSs (gNB) on rooftops, corresponding to an inter-site distance (ISD) of approximately , typical for urban microcellular evaluations. Terrestrial UEs are randomly placed outside buildings, and ray tracing was performed using Wireless Insite [Remcom], which has recently been proven fairly accurate by conducting real-world measurements in the mmWave bands, as reported in [de2023ray, remcommpresnt2017modeling].
As mentioned in the Introduction, a critical factor in the performance of mmWave systems is the susceptibility of signals to blockage [Rappaport2014-mmwbook, slezak2018empirical, slezak2018understanding, maccartney2017rapid, zekri2020analysis]. To model the effect of blockage, we consider two scenarios:
No blockage; and
Blockage modeled with the 3GPP Blockage Model B [3GPP38901] with random human blockers.
Fig. 2c shows a coverage map of the unblocked wideband SNR for a single BS deployment as an example. As expected, we see that coverage is greatly reduced as frequency increases.
@@
For both the unblocked and blocked cases, we compute the wideband SNR for each UE and BS and assume the UE is served by the BS with the strongest unblocked SNR. Then, given an , we assume the achieved rate (goodput) follows a standard realistic model [mogensen2007lte]:
@@
where is the bandwidth, is a system bandwidth loss factor to account for overhead and receiver imperfections, and is a maximum spectral efficiency. Following [mogensen2007lte], we adopt for and for .
@@
Fig. 3 shows the resulting SNR distribution for frequencies 6, 12, 18 and 24 at all outdoor UE locations in the study area without blockage. As can be seen in Fig. 3, UEs at lower frequencies experience uniformly better SNRs than higher frequencies due to favorable propagation and reduced noise power from using a smaller bandwidth compared to higher frequencies. Fig. 4a shows the corresponding rate distribution. We observe that, even though the SNRs are lower, as expected, the use of higher frequencies (18 and 24 ) ensures superior data rate due to the wider bandwidth. Note that from (1), there is a maximum rate of which corresponds to , , and at frequencies , , and , respectively. Hence, higher peak rates can be achieved at higher frequencies, where more bandwidth is available. However, UEs at the cell edge (for example, UEs up to the bottom 10% percentile of the CDF) suffer a significantly worse rate in the higher frequency bands (18 and 24 ) compared to UEs at lower frequencies (6 and 12 ). This property is also expected, as propagation is much less favorable for UEs at the cell edge at higher frequencies.
Fig. 4b shows the rate distribution with blockage. We observe that the cell edge rates at the high frequencies drop sharply. Indeed, UEs in and show significantly worse performance approximately of the time, compared to UEs in and . Thus, when blockers, such as humans or vehicles, surround an UE, as would occur commonly in an urban scenario, the high-frequency coverage suffers significantly, while the low-frequency coverage is significantly more uniform.
@@
The curves labeled “best choice” in Fig. 4a and Fig. 4b correspond to the rates for the UEs that select the best BS and frequency. As expected, the best choice is uniformly better than any individual frequency. Indeed, the best choice achieves the same peak rates as the UEs in the high-frequency bands, along with the improved cell edge rates of the low frequencies. These results motivate wideband cellular systems that adaptively select across a range of bands. Such systems can obtain the bandwidth benefits at high frequencies while providing robustness and resistance to blockage at low frequencies.
3.3 Penetration Loss
@@
3.4 Multi-frequency Indoor Capacity Gains
To understand the effect of penetration loss on capacity, we re-run the simulation with users randomly placed indoors. Note that the 3D models used in the ray tracing have approximate material classifications for the facades of each building. Using the penetration loss models from Eq. 2 combined with the exterior wall classification, Fig. 6 plots the estimated CDFs of the SNR and data rate for indoor UEs. Fig. 6a shows that the SNR for indoor UEs is considerably reduced at higher frequencies due to high penetration loss, as mentioned in the previous section. In this setting, concrete is a dominant exterior wall material and significantly reduces the signal penetration – see Fig. 5. Indeed, Fig. 6b shows that, for roughly 65% of the users, the data rates at lower frequencies are higher than at higher frequencies. These results indicate a further potential gain of wideband cellular systems in FR3: the lower frequencies can provide valuable indoor coverage behind materials such as concrete while the higher frequencies can opportunistically offer high capacity for outdoor users and selected indoor users with minimal blockage (e.g., indoor users next to non-IRR glass windows).
3.5 Effects of Interference